# Fast Generation of Feasible Trajectories in Direct Optimal Control

### v1

@@ III-C Line Search and Step Acceptance
Due to the use of DDP, globalization is only required for the objective function which is once differentiable due to the choice of the squared 2-norm. A prominent feature of this choice is that the Maratos effect is avoided [17]. The backward Riccati recursion on problem (III-A) yields the gain matrices $K_{i}$ and the feedforward matrices $k_{i}$. The forward sweep, given by (6a), (6b), and (7), includes the line search parameter $\alpha\in(0,1]$. An Armijo condition decides upon step acceptance. Therefore, the predicted reduction of the quadratic model of the objective function $f$ at iterate $(\mathbf{x}^{(k)},\mathbf{u}^{(k)})$ using problem (III-A) and its solution $(\mathbf{\delta x},\mathbf{\delta u})$ is defined by
$\begin{split}m_{f}^{(k)}(\mathbf{\delta x},\mathbf{\delta u})=&-(\nabla f^{(k)})^{\top}(\mathbf{\delta x},\mathbf{\delta u})\\ &-\frac{1}{2}(\mathbf{\delta x},\mathbf{\delta u})^{\top}(H^{\mathrm{GGN}})^{(k)}(\mathbf{\delta x},\mathbf{\delta u}).\end{split}$ (10)
@@
Note that $m_{f}^{(k)}(\mathbf{\delta x},\mathbf{\delta u})=-f^{*}_{\mathrm{QP}}(\mathbf{x}^{(k)},\mathbf{u}^{(k)})$. Given the current iterate $(\mathbf{x}^{(k)},\mathbf{u}^{(k)})$, the trial iterate for a given $\alpha$ is denoted by $(\mathbf{\hat{x}}(\alpha),\mathbf{\hat{u}}(\alpha))$. The Armijo condition is defined by
@@
for $\eta\in(0,1)$. If the current step size is not acceptable for the Armijo condition (11), the step size is reduced by a factor $\tfrac{1}{2}$, i.e., $\alpha\leftarrow\tfrac{1}{2}\alpha$.
III-D Initialization and Termination
As mentioned in Section III-A, DDP needs to be initialized at a dynamically feasible point. In case such a guess is not provided, prior to starting FP-DDP one DDP backward-forward simulation can generate a dynamically feasible trajectory.
The algorithm stops, if a feasible point is found, i.e., the objective function is below a given tolerance $F(\mathbf{u})\leq\varepsilon_{\mathrm{F}}$ or if a stationary point was found, i.e., $\|\nabla F(\mathbf{u})\|_{\infty}\leq\varepsilon_{\mathrm{S}}$.
Parameter : $\eta,\alpha_{\min},\varepsilon_{\mathrm{F}},\varepsilon_{\mathrm{S}},\mu_{\min}\in(0,1),\lambda,\tau>0;$
@@ III-D Initialization and Termination
1 Find dynamically feasible initial guess $(\mathbf{x}^{(0)},\mathbf{u}^{(0)})$ for $k=0,1,2,\ldots$ do
2       if $F(\mathbf{u})\leq\varepsilon_{\mathrm{F}}$ or $\|\nabla F(\mathbf{u})\|_{\infty}\leq\varepsilon_{\mathrm{\mathrm{S}}}$ then
3            break
4      Setup QP (III-A) and calculate $K_{i},k_{i}$ from (3), (2) $\alpha^{(k)}\leftarrow 1.0$, calculate $m_{f}^{(k)}(\mathbf{\delta x},\mathbf{\delta u})$ with (8) while $\mathrm{True}$ do
5             if $\alpha^{(k)}<\alpha_{\min}$ then
6                   $\mu^{(k+1)}\leftarrow\tau\mu^{(k)}$ break
7            Calculate $(\mathbf{\hat{x}}^{(k)}(\alpha),\mathbf{\hat{u}}^{(k)}(\alpha))$ from (6a), (6b),(7) if Sufficient decrease (11) is satisfied then
8                   $(\mathbf{x}^{(k+1)},\mathbf{u}^{(k+1)})\leftarrow(\mathbf{\hat{x}}^{(k)}(\alpha),\mathbf{\hat{u}}^{(k)}(\alpha))$ Set $\mu^{(k+1)}$ depending on $\alpha^{(k)}$ break
9             else
10                  $\alpha^{(k)}\leftarrow\frac{1}{2}\alpha^{(k)}$
11
12
@@ 10                  $\alpha^{(k)}\leftarrow\frac{1}{2}\alpha^{(k)}$
IV CONVERGENCE ANALYSIS
In this section a global convergence result for FP-DDP is presented. Local convergence follows from the results in [2] and [3]. The first global convergence proof for DDP was given in [23]. The derivation below is adapted to FP-DDP. Before giving a proof, several auxiliary lemmas are introduced.
@@ IV CONVERGENCE ANALYSIS
Let a trajectory $(\bar{\mathbf{x}},\bar{\mathbf{u}})$ be given that satisfies the dynamics (II-B), (II-B). Then, the following statements hold:
@@
    If $(\bar{\mathbf{x}},\bar{\mathbf{u}})$ is not optimal, then the Hessian matrix constructed by Algorithm (1) is positive definite.
@@
    The optimal solution of (III-A) is $(\mathbf{\delta x},\mathbf{\delta u})=0$ if and only if $(\mathbf{\bar{x}},\mathbf{\bar{u}})$ is a KKT point of (II-B).
@@
    The predicted reduction of (III-A) is zero, $m_{f}(\mathbf{\delta x},\mathbf{\delta u})=0$, if and only if $(\mathbf{\delta x},\mathbf{\delta u})=0$. Otherwise $m_{f}(\mathbf{\delta x},\mathbf{\delta u})>0$.
@@
Since $(\bar{\mathbf{x}},\bar{\mathbf{u}})$ is not optimal, $f(\bar{\mathbf{x}},\bar{\mathbf{u}})$ is non-zero. Thus, the regularization term in (9) is positive definite which shows (i). Result (ii) directly follows from comparison of the KKT conditions of (II-B) and (III-A) with $(\mathbf{\delta x},\mathbf{\delta u})=0$. For (iii), it is noted that all iterates of Algorithm 1 are feasible for the dynamics. Therefore, QP (III-A) has a feasible direction with $(\mathbf{\delta x},\mathbf{\delta u})=0$ for which the objective function value of (III-A) is 00. Noting that the Hessian matrix is positive definite implies that there exists a unique solution to (III-A). Recalling that $m_{f}((\mathbf{\delta x},\mathbf{\delta u}))=-f_{\mathrm{QP}}^{*}(\mathbf{x},\mathbf{u})$ shows that $m_{f}((\mathbf{\delta x},\mathbf{\delta u}))=0$ if $(\mathbf{\delta x},\mathbf{\delta u})=0$ and otherwise $m_{f}((\mathbf{\delta x},\mathbf{\delta u}))>0$. ∎
@@
For the DDP step it holds that
@@
The lemma can be proved by induction and comparing the derivative of $(\hat{\mathbf{x}}(\alpha),\hat{\mathbf{u}}(\alpha))$ at $\alpha$ with $(\mathbf{\delta x},\mathbf{\delta u})$ obtained from (6a)-(6c). ∎
Lemma 3 ([23]).
For the objective function with respect to $\mathbf{u}$, i.e., $F$, it holds
@@
Let a trajectory $(\bar{\mathbf{x}},\bar{\mathbf{u}})$ be given that satisfies the dynamics (II-B) and that is not optimal and let $\eta\in(0,1)$. then there exists an $\alpha\in(0,1]$ such that the sufficient decrease condition (11) is satisfied.
@@
Let the trial iterate be denoted by $(\hat{\mathbf{x}}(\alpha),\hat{\mathbf{u}}(\alpha))$. Noting that $F(\hat{\mathbf{u}}(\alpha))=f(\hat{\mathbf{x}}(\alpha),\hat{\mathbf{u}}(\alpha))$, $F(\bar{\mathbf{u}})=f(\bar{\mathbf{x}},\bar{\mathbf{u}})$, and $-f^{*}_{\mathrm{QP}}(\bar{\mathbf{x}},\bar{\mathbf{u}})=m_{f}(\mathbf{\delta x},\mathbf{\delta u})$ and using Lemma 3, it follows
$f(\bar{\mathbf{x}},\bar{\mathbf{u}})-f(\hat{\mathbf{x}}(\alpha),\hat{\mathbf{u}}(\alpha))\geq\alpha m_{f}(\mathbf{\delta x},\mathbf{\delta u})-\bar{\xi}\alpha^{2}.$
for $\bar{\xi}>0$. Choosing $\alpha\leq((1-\eta)m_{f}(\mathbf{\delta x},\mathbf{\delta u}))/\bar{\xi}$ yields the claim. ∎
@@
Suppose that Algorithm 1 creates a sequence of iterates $\{(\mathbf{x}^{(k)},\mathbf{u}^{(k)})\}_{k\in\mathbb{N}}$. Suppose that $\nabla f$ is Lipschitz continuous and that $H^{\mathrm{GGN}}$ is uniformly bounded for all $k$. Then, any accumulation point $(\mathbf{x}^{*},\mathbf{u}^{*})$ of the sequence is a stationary point for (II-B).
@@
For all iterations $k$ the sufficient decrease condition (11) is satisfied. Summing for $k=0$ to $N$ and using a telescopic sum argument yields
$\eta\sum_{k=0}^{N}\alpha^{(k)}m_{f}^{(k)}(\mathbf{\delta x}^{(k)},\mathbf{\delta x}^{(k)})\leq f^{(0)}-f^{(N)}\leq f^{(0)}$
@@
$\sum_{k=0}^{\infty}\alpha^{(k)}m_{f}^{(k)}(\mathbf{\delta x}^{(k)},\mathbf{\delta u}^{(k)})<\infty,$ (12)
implying $\alpha^{(k)}m_{f}^{(k)}(\mathbf{\delta x}^{(k)},\mathbf{\delta u}^{(k)})\to 0$. To finish the proof it needs to be shown that $\alpha^{(k)}$ is bounded away from zero for $k\to\infty$. For a contradiction, assume that the predicted reduction $m_{f}^{(k)}$ does not converge to 0, i.e., $(\mathbf{\delta x}^{(k)},\mathbf{\delta u}^{(k)})\not\to 0$. From this follows $\alpha^{(k)}\to 0$.
If $\alpha^{(k)}=1$, there is nothing to show. So, assume that for $\alpha^{(k)}<1$ the Armijo condition (11) is satisfied and dropping the index $k$, then for $2\alpha$, (11) is not satisfied. Equation (11) being not satisfied is equivalent to
$F(\mathbf{\hat{u}}(2\alpha))-F(\mathbf{\bar{u}})>2\eta\alpha\nabla F(\mathbf{\bar{u}})^{\top}\delta\mathbf{u}.$ (13)
@@
From here, knowing that $F$ is once continuously differentiable with Lipschitz gradients and bounded Hessian a standard unconstrained optimization proof from, e.g., [7, Theorem 9.1] can be applied to bound $\alpha$ from below. The DDP step needs to be expressed in terms of $\delta\mathbf{u}$ using Lemma 2. Then, it can be shown that $\alpha>M\|\delta\mathbf{u}\|^{2}$ for $M>0$. From this follows that $\alpha\not\to 0$ which is a contradiction to (12). ∎
V Simulation Results
@@
This section presents simulation results of FP-DDP that demonstrate its efficiency. Two test problems are considered: (i) a fixed-time P2P motion of an unstable system; (ii) a free-time P2P motion of a cart pendulum system with a circular obstacle.
V-A Implementation
@@
The OCPs are implemented using the software rockit [12] which is a rapid prototyping OCP toolbox built on top of CasADi [1]. The HPIPM Python interface [10] is used to generate the Riccati matrices. A prototypical python implementation of the FP-DDP algorithm is available at github.com/david0oo/fp_ddp_python. The parameters of FP-DDP were chosen as $\eta=10^{-6}$, $\alpha_{\min}=10^{-17}$, $\mu_{\min}=10^{-16}$, $\mu^{(0)}=1.0$, $\lambda=5$, $\tau=10$, $\varepsilon_{\mathrm{F}}=10^{-12}$, and $\varepsilon_{\mathrm{S}}=10^{-8}$. In example (ii), the IPOPT default settings were used. For the truncated Newton method of scipy.optimize.minimize, $\mathrm{tol}=10^{-8}$ was used. All simulations were carried out using an Intel core i7-10810U CPU.
V-B Fixed-Time P2P Motion of Unstable System

### v2

@@ 3.3 Line Search and Step Acceptance
Due to the use of DDP, globalization is only required for the objective function which is once differentiable due to the choice of the squared 2-norm. A prominent feature of this choice is that the Maratos effect is avoided [17]. The backward Riccati recursion on problem (3.1) yields the gain matrices $K_{i}$ and the feedforward matrices $k_{i}$. The forward sweep, given by (7a) and (8), includes the line search parameter $\alpha\in(0,1]$. An Armijo condition decides upon step acceptance. Therefore, the predicted reduction of the quadratic model of $f$ at iterate $(\mathbf{x}^{(k)},\mathbf{u}^{(k)})$ using problem (3.1) and its solution $(\mathbf{\delta x},\mathbf{\delta u})$ is defined by
[…]
@@
Note that $m_{f}^{(k)}(\mathbf{\delta x},\mathbf{\delta u})=-f^{*}_{\mathrm{QP}}(\mathbf{x}^{(k)},\mathbf{u}^{(k)})$. Given the iterate $(\mathbf{x}^{(k)},\mathbf{u}^{(k)})$, the trial iterate for a given $\alpha$ is denoted by $(\mathbf{\hat{x}}(\alpha),\mathbf{\hat{u}}(\alpha))$. The Armijo condition is defined by
@@
for $\eta\in(0,1)$. If the current step size is not acceptable for (12), the step size is reduced by a factor $\tfrac{1}{2}$, i.e., $\alpha\leftarrow\tfrac{1}{2}\alpha$.
3.4 Initialization and Termination
As mentioned in Section 3.1, DDP requires a dynamically feasible initialization. In case such a guess is not provided, prior to starting FP-DDP, a Riccati backward sweep (3) - (6) is performed based on the infeasible initial guess and a nonlinear forward simulation (7a) and (8) yields the dynamically feasible initialization. The algorithm stops, if a feasible trajectory is found, i.e., the objective function is below a given tolerance $F(\mathbf{u})\leq\varepsilon_{\mathrm{F}}$ or if a stationary point was found, i.e., $\|\nabla F(\mathbf{u})\|_{\infty}\leq\varepsilon_{\mathrm{S}}$.
Parameter : $\eta,\alpha_{\min},\varepsilon_{\mathrm{F}},\varepsilon_{\mathrm{S}},\mu_{\min}\in(0,1),\lambda>0;$
@@ 3.4 Initialization and Termination
1 Find dynamically feasible initial guess $(\mathbf{x}^{(0)},\mathbf{u}^{(0)})$
2 for $k=0,1,2,\ldots$ do
3       if $F(\mathbf{u})\leq\varepsilon_{\mathrm{F}}$ or $\|\nabla F(\mathbf{u})\|_{\infty}\leq\varepsilon_{\mathrm{\mathrm{S}}}$ then break
4       Setup QP (3.1) and calculate $K_{i},k_{i}$ from (4), (3)
5       $\alpha^{(k)}\leftarrow 1.0$, calculate $m_{f}^{(k)}(\mathbf{\delta x},\mathbf{\delta u})$ with (9)
6       while $\mathrm{True}$ do
7             if $\alpha^{(k)}<\alpha_{\min}$ then  $\mu^{(k+1)}\leftarrow\lambda\mu^{(k)}$; break
8             Calculate $(\mathbf{\hat{x}}^{(k)}(\alpha),\mathbf{\hat{u}}^{(k)}(\alpha))$ from (7a), (8)
9             if Sufficient decrease (12) is satisfied then
10                   $(\mathbf{x}^{(k+1)},\mathbf{u}^{(k+1)})\leftarrow(\mathbf{\hat{x}}^{(k)}(\alpha),\mathbf{\hat{u}}^{(k)}(\alpha))$
11                   Set $\mu^{(k+1)}$ depending on $\alpha^{(k)}$; break
12             else $\alpha^{(k)}\leftarrow\frac{1}{2}\alpha^{(k)}$
13
14
@@ 11                   Set $\mu^{(k+1)}$ depending on $\alpha^{(k)}$; break
4 Convergence Analysis
This section presents a global convergence result for FP-DDP. Local convergence follows from the results in [2] and [3]. Firstly, several auxiliary lemmas are introduced.
@@ 4 Convergence Analysis
Let a trajectory $(\bar{\mathbf{x}},\bar{\mathbf{u}})$ be given that satisfies the dynamics (2.2). Then, the following statements hold:
@@
    If $(\bar{\mathbf{x}},\bar{\mathbf{u}})$ is not optimal for (2.2), then the Hessian matrix constructed by Algorithm 1 is positive definite.
@@
    The optimal solution of (3.1) is $(\mathbf{\delta x},\mathbf{\delta u})=0$ if and only if $(\mathbf{\bar{x}},\mathbf{\bar{u}})$ is a KKT point of (2.2).
@@
    The predicted reduction of (3.1) is zero, $m_{f}(\mathbf{\delta x},\mathbf{\delta u})=0$, iff $(\mathbf{\delta x},\mathbf{\delta u})=0$. Otherwise $m_{f}(\mathbf{\delta x},\mathbf{\delta u})>0$.
@@
Since $(\bar{\mathbf{x}},\bar{\mathbf{u}})$ is not optimal, $f(\bar{\mathbf{x}},\bar{\mathbf{u}})$ is non-zero, i.e., the regularization term in (10) is positive. Thus, referring to Sec. 3.2 yields (i). Result (ii) directly follows from comparing the KKT conditions of (2.2) and (3.1) with $(\mathbf{\delta x},\mathbf{\delta u})\!=\!0$. For (iii): Since all iterates of Algorithm 1 are dynamically feasible, QP (3.1) has a feasible direction with $(\mathbf{\delta x},\mathbf{\delta u})=0$ for which the objective function value of (3.1) is 00. The positive definiteness of the Hessian implies that there exists a unique solution to (3.1). Recalling that $m_{f}((\mathbf{\delta x},\mathbf{\delta u}))=-f_{\mathrm{QP}}^{*}(\bar{\mathbf{x}},\bar{\mathbf{u}})$ shows that $m_{f}((\mathbf{\delta x},\mathbf{\delta u}))=0$ if $(\mathbf{\delta x},\mathbf{\delta u})=0$ and otherwise $m_{f}((\mathbf{\delta x},\mathbf{\delta u}))>0$. ∎
@@
For the DDP step (7a), (8) it holds that
@@
The lemma can be proved by induction and comparing the derivative of $(\hat{\mathbf{x}}(\alpha),\hat{\mathbf{u}}(\alpha))$ at $\alpha$ with $(\mathbf{\delta x},\mathbf{\delta u})$ obtained from (7). ∎
Lemma 3 (from [24]).
For the objective function $F$ with eliminated states, it holds
@@
Let a dynamically feasible trajectory $(\bar{\mathbf{x}},\bar{\mathbf{u}})$ be given that is not optimal and let $\eta\in(0,1)$. Then there exists an $\alpha\in(0,1]$ such that the sufficient decrease condition (12) is satisfied.
@@
Let the trial iterate be denoted by $(\hat{\mathbf{x}}(\alpha),\hat{\mathbf{u}}(\alpha))$. Noting that $F(\hat{\mathbf{u}}(\alpha))=f(\hat{\mathbf{x}}(\alpha),\hat{\mathbf{u}}(\alpha))$, $F(\bar{\mathbf{u}})=f(\bar{\mathbf{x}},\bar{\mathbf{u}})$, and $-f^{*}_{\mathrm{QP}}(\bar{\mathbf{x}},\bar{\mathbf{u}})=m_{f}(\mathbf{\delta x},\mathbf{\delta u})$ and using Lemma 3, it follows $f(\bar{\mathbf{x}},\bar{\mathbf{u}})-f(\hat{\mathbf{x}}(\alpha),\hat{\mathbf{u}}(\alpha))\geq\alpha m_{f}(\mathbf{\delta x},\mathbf{\delta u})-\bar{\xi}\alpha^{2}$ for $\bar{\xi}>0$, i.e., $\alpha\leq((1-\eta)m_{f}(\mathbf{\delta x},\mathbf{\delta u}))/\bar{\xi}$ yields the claim. ∎
@@
Suppose that $\nabla f$ is Lipschitz continuous and that $H^{(k)}$ is uniformly bounded for all $k$. Then, Algorithm 1 creates a sequence of iterates $\{(\mathbf{x}^{(k)},\mathbf{u}^{(k)})\}_{k\in\mathbb{N}}$ and any accumulation point $(\mathbf{x}^{*},\mathbf{u}^{*})$ of the sequence is a stationary point for (2.2).
@@
From Lemma 4 follows that Algorithm 1 is well-defined, i.e., a sequence $\{(\mathbf{x}^{(k)},\mathbf{u}^{(k)})\}_{k\in\mathbb{N}}$ is found and that the sufficient decrease condition (12) is satisfied in every iteration $k$. Summing for $k=0$ to $N$ and using a telescopic sum argument yields
@@
$\sum_{k=0}^{\infty}\alpha^{(k)}m_{f}^{(k)}(\mathbf{\delta x}^{(k)},\mathbf{\delta u}^{(k)})<\infty,$
@@
implying $\alpha^{(k)}m_{f}^{(k)}(\mathbf{\delta x}^{(k)},\mathbf{\delta u}^{(k)})\to 0$ since Lemma 1 (iii) guarantees non-negativity of the summands. To finish the proof it needs to be shown that $\alpha^{(k)}$ is bounded away from zero for $k\to\infty$. For a contradiction, assume that the predicted reduction $m_{f}^{(k)}$ does not converge to 0, i.e., $(\mathbf{\delta x}^{(k)},\mathbf{\delta u}^{(k)})\not\to 0$. This implies $\alpha^{(k)}\to 0$. If $\alpha^{(k)}=1$, there is nothing to show. So, assume that for $\alpha^{(k)}<1$ the Armijo condition (12) is satisfied and dropping the index $k$, then for $2\alpha$, (12) is not satisfied. Equation (12) being not satisfied is equivalent to $F(\mathbf{\hat{u}}(2\alpha))-F(\mathbf{\bar{u}})>2\eta\alpha\nabla F(\mathbf{\bar{u}})^{\top}\delta\mathbf{u}$. From here, knowing that $F$ is once continuously differentiable with Lipschitz gradients and bounded Hessian a standard unconstrained optimization proof, e.g., [6, Thm 9.1], can be applied to bound $\alpha$ from below. The DDP step needs to be expressed in terms of $\delta\mathbf{u}$ using Lemma 2. Then, it can be shown that $\alpha>M\|\delta\mathbf{u}\|^{2}$ for $M>0$. From this, it follows that $\alpha\not\to 0$ which is a contradiction to (13). As a consequence, $m_{f}^{(k)}(\mathbf{\delta x}^{(k)},\mathbf{\delta u}^{(k)})\to 0$ and convergence follows from Lemma 1. ∎
Theorem 1 guarantees convergence of FP-DDP to a KKT point, but in general, it does not guarantee that $f(\mathbf{x},\mathbf{u})=0$ at the solution. If the OCP (1) is convex, FOCP (2.2) is also convex. If therefore, the original OCP has feasible points, all local solutions of (2.2) are global solutions and fulfill $f(\mathbf{x},\mathbf{u})=0$. In this case, FP-DDP is guaranteed to converge to a feasible point. For nonconvex problems, this cannot be guaranteed since many local solutions can exist. If the optimal objective function value is nonzero, this indicates that the problem might be locally infeasible [22]. In the simulation examples presented next, this case never occurred. FP-DDP always found a feasible solution.
5 Simulation Results
@@
This section demonstrates the capabilities of FP-DDP on two test problems, (i) a fixed-time P2P motion of an unstable system, (ii) a free-time P2P motion of a cart pendulum system with a circular obstacle.
5.1 Implementation
@@
The OCPs are formulated using the rapid prototyping toolbox rockit [10] built on top of CasADi [1]. The FP-DDP algorithm is implemented within acados [19] utilizing the Riccati recursion from HPIPM [9]. The code to reproduce the results and a prototypical python implementation of FP-DDP are publicly available [12]. The parameters of FP-DDP were chosen as $\eta=10^{-6}$, $\alpha_{\min}=10^{-17}$, $\mu_{\min}=10^{-16}$, $\mu^{(0)}=10^{-3}$, $\lambda=5$, $\varepsilon_{\mathrm{F}}=10^{-12}$, and $\varepsilon_{\mathrm{S}}=10^{-8}$. In example (ii), the IPOPT default settings were used. For the truncated Newton method of scipy.optimize.minimize, $\mathrm{tol}=10^{-8}$ was used. All simulations were carried out using an i7-10810U CPU.
5.2 Fixed-Time P2P Motion of Unstable System